# Low-Dimensional Federated Knowledge Graph Embedding via Knowledge Distillation

### v1

@@ 1. Introduction
  • We assess the efficacy of FedKD across three datasets through comprehensive experiments. The findings indicate a notable reduction in the embedding dimension of FKGE with with only marginal (even no) decrease in performance.
@@ 2.1. Federated Knowledge Graph Embedding
Federated Learning (FL) represents a novel distributed paradigm in machine learning that enables multiple clients to collaboratively train high-performing models, simultaneously preserving data privacy and security (Li et al., 2020, 2021b; T Dinh et al., 2020; Chen et al., 2022a; Zhang et al., 2024d). Based on the FL paradigm, federated knowledge graph embedding aims to facilitate collaborative learning of entity and relation embeddings from multi-source KGs distributed across multiple clients, while ensuring raw triplets are kept confidential and not shared among participants.
FedE (Chen et al., 2021), inspired by FedAvg (McMahan et al., 2017), serves as a pioneering model where a server coordinates clients’ local embedding learning by iteratively aggregating their entity embeddings and distributing the aggregated result for the subsequent round of local training. Inspired by Moon (Li et al., 2021a), FedEC introduces embedding contrastive learning into FedE to further improve embedding performance. However, both FedE and FedEC generate a single set of global consensus embeddings for all clients, which may not fully accommodate data heterogeneity. As a result, the embeddings trained to fit an “average KG” might not generalize well to KGs that exhibit high heterogeneity with the global training dataset. To address it, several personalized FKGE methods are proposed. PFedEG proposed the personalized embedding aggregation mechanism on the server based on a client-wise relation graph (Zhang et al., 2024c). FedLu introduces mutual knowledge distillation to facilitate the transfer of information from local entity embeddings to global entity embeddings, followed by the incorporation of knowledge gleaned from the global entity embeddings (Zhu et al., 2023). In contrast to the aforementioned methods that predominantly handle scenarios where clients possess partially shared entities but mutually exclusive relations, FedR (Zhang et al., 2022) specifically addresses cases where clients not only share entities but also share relations. In this framework, all clients receive identical embeddings of shared relations from the server and subsequently engage in local embedding learning using their respective local triples together with the shared relation embeddings.
Unlike the aforementioned methods that rely on a client-server architecture, FKGE (Peng et al., 2021) facilitates collaborative embedding learning among clients in a peer-to-peer manner. Drawing inspiration from MUSE (Conneau et al., 2017), FKGE adopts a Generative Adversarial Network (GAN) framework (Goodfellow et al., 2020) to harmonize the embeddings of shared entities and relations within a pair of knowledge graphs. Besides, FedS (Zhang et al., 2024b) focus on reducing the communication cost of FKGE methods based on the client-server architecture, by introducing entity-wise embedding sparsification strategy.
@@ 3. Preliminaries
FedE is the first FKGE method which mainly includes two parts: client update and server update. We take the communication round $t$ as an example to explain them, considering the different rounds share the same process. For server updates, the server first collects the updated entity embeddings at round $t-1$ from each client. Subsequently, the server performs an average embedding aggregation for each entity. For client updates, each client initially receives the aggregated embeddings from the server and replaces their local entity embeddings with these updated versions. Following this, each client proceeds with local embedding learning using a KGE method. The objective of local embedding learning is scoring positive triples higher than negative ones based on the score function of the chosen KGE method.
FKGE Compression based on Knowledge Distillation (FKGE-CD) aims at reducing the embedding dimension for a FKGE method using knowledge distillation techniques, and simultaneously avoiding decrease embedding performance significantly. In this paper, we further assume that the FKGE-CD training process should follow the existing client and server communication mechanisms, and it should only intervene in the client local learning process. Other scenarios are reserved for future research.
@@ 4.2. Client Local Training with Knowledge Distillation
(4) ($\mathcal{L}^{S}_{(h,r,t)}=KL(\mathcal{P}^{stu}_{(h,r,\{t,t^{\prime}_{i}\})},\mathcal{P}^{tea}_{(h,r,\{t,t^{\prime}_{i}\})})$
where $\{t,t^{\prime}_{i}\}$ ($i\in[1,n]$) denotes the tail entity set of triple $(h,r,t)$ and its negative samples; $KL(\cdot)$ denotes the Kullback-Leiber Divergence; $\mathcal{P}^{stu}_{(h,r,\{t,t^{\prime}_{i}\})}$ and $\mathcal{P}^{tea}_{(h,r,\{t,t^{\prime}_{i}\})}$ are the score distributions of the student and the teacher model about triple $(h,r,t)$ and its negative samples, respectively. For student model, $\mathcal{P}^{stu}_{(h,r,\{t,t^{\prime}_{i}\})}$ is composed by the score probability $\mathcal{P}^{stu}_{(h,r,\hat{t})}$ ($\hat{t}\in\{t,t^{\prime}_{i}\}$) of triple $(h,r,t)$ and its negative samples. $\mathcal{P}^{stu}_{(h,r,\hat{t})}$ is computed as follows:
(5) $\mathcal{P}^{stu}_{(h,r,\hat{t})}=\frac{exp(S^{(h,r,\hat{t})}_{stu}/\tau)}{\sum_{\hat{t}\in\{t,t^{\prime}_{i}\}}exp(S^{(h,r,\hat{t})}_{stu}/\tau)}$

### v2

@@ 1. Introduction
  • We assess the efficacy of FedKD across three datasets through comprehensive experiments. The findings indicate a notable reduction in the embedding dimension of FKGE with only marginal (even no) decrease in performance.
@@ 2.1. Federated Knowledge Graph Embedding
Federated Learning (FL) represents a novel distributed paradigm in machine learning that enables multiple clients to collaboratively train high-performing models, simultaneously preserving data privacy and security (Li et al., 2020, 2021b; T Dinh et al., 2020; Chen et al., 2022a; Zhang et al., 2024b). Based on the FL paradigm, federated knowledge graph embedding aims to facilitate collaborative learning of entity and relation embeddings from multi-source KGs distributed across multiple clients, while ensuring raw triplets are kept confidential and not shared among participants.
FedE (Chen et al., 2021), inspired by FedAvg (McMahan et al., 2017), serves as a pioneering model where a server coordinates clients’ local embedding learning by iteratively aggregating their entity embeddings and distributing the aggregated result for the subsequent round of local training. Inspired by Moon (Li et al., 2021a), FedEC introduces embedding contrastive learning into FedE to further improve embedding performance. However, both FedE and FedEC generate a single set of global consensus embeddings for all clients, which may not fully accommodate data heterogeneity. As a result, the embeddings trained to fit an “average KG” might not generalize well to KGs that exhibit high heterogeneity with the global training dataset. To address it, several personalized FKGE methods are proposed. PFedEG proposed the personalized embedding aggregation mechanism on the server based on a client-wise relation graph (Zhang et al., 2025b). FedLu introduces mutual knowledge distillation to facilitate the transfer of information from local entity embeddings to global entity embeddings, followed by the incorporation of knowledge gleaned from the global entity embeddings (Zhu et al., 2023). In contrast to the aforementioned methods that predominantly handle scenarios where clients possess partially shared entities but mutually exclusive relations, FedR (Zhang et al., 2022) specifically addresses cases where clients not only share entities but also share relations. In this framework, all clients receive identical embeddings of shared relations from the server and subsequently engage in local embedding learning using their respective local triples together with the shared relation embeddings.
Unlike the aforementioned methods that rely on a client-server architecture, FKGE (Peng et al., 2021) facilitates collaborative embedding learning among clients in a peer-to-peer manner. Drawing inspiration from MUSE (Conneau et al., 2017), FKGE adopts a Generative Adversarial Network (GAN) framework (Goodfellow et al., 2020) to harmonize the embeddings of shared entities and relations within a pair of knowledge graphs. Besides, FedS (Zhang et al., 2025a) focus on reducing the communication cost of FKGE methods based on the client-server architecture, by introducing entity-wise embedding sparsification strategy.
@@ 3. Preliminaries
FedE is the first FKGE method which mainly includes two parts: client update and server update. We take the communication round $t$ as an example to explain them, considering the different rounds share the same process. For server updates, the server first collects the updated entity embeddings at round t1t-1 from each client. Subsequently, the server performs an average embedding aggregation for each entity. For client updates, each client initially receives the aggregated embeddings from the server and replaces their local entity embeddings with these updated versions. Following this, each client proceeds with local embedding learning using a KGE method. The objective of local embedding learning is scoring positive triples higher than negative ones based on the score function of the chosen KGE method.
FKGE Compression based on Knowledge Distillation (FKGE-CKD) aims at reducing the embedding dimension for a FKGE method using knowledge distillation techniques, and simultaneously avoiding decrease embedding performance significantly. In this paper, we further assume that the FKGE-CKD training process should follow the existing client and server communication mechanisms, and it should only intervene in the client local learning process. Other scenarios are reserved for future research.
@@ 4.2. Client Local Training with Knowledge Distillation
[…]